Circular No. 8051

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002kg IN NGC 2403 Further to IAUC 8048, M. Schwartz and W. Li report the LOTOSS

discovery of a supernova on unfiltered images taken with the Tenagra II telescope and KAIT. The object was at mag 19 ± 0.3, and possibly showed a brightening trend, from 2002 Oct. 26 to 2003 Jan. 1 UT; SN 2002kg is present in Tenagra II images taken on 2002 Oct. 30, Nov. 14, 22, 24, 25, Dec. 8, and 2003 Jan. 1, and in KAIT images taken on 2002 Oct. 26, Nov. 1, and Dec. 2. The object is located at α = 7^{h}37^{m}01^{s}.83, δ = +65^{o}34′29′′.3 (equinox 2000.0), which is 71′′.7 east and 96′′.0 south of the nucleus of NGC 2403. A KAIT image of the same field taken on 1998 Nov. 13 showed nothing at this position (limiting mag ∼ 19.5). A. V. Filippenko and R. Chornock, University of California, Berkeley, write: “Inspection of CCD spectra (range 310–1000 nm) obtained on Jan. 6 UT with the Keck I 10-m telescope (+ LRIS) shows that SN 2002kg is of type IIn, specifically the variety similar to SN 1997bs (Van Dyk et al. 2000, PASP 112, 1532). The narrow Balmer emission lines (unresolved; FWHM < 500 km/s) have a broader base with FWHM roughly 2500 km/s (measured from Hβ), and there is weak Fe II emission at 450–460 and 515–535 nm. Prominent, unresolved [N II] 654.8- and 658.4-nm emission lines are visible, yet other forbidden lines are weak or absent, suggesting a possible nitrogen overabundance in the circumstellar gas. Similar recent examples of such objects (but without the [N II] emission) include SN 1999bw (IAUC 7152), SN 2000ch (IAUC 7421), and SN 2001ac (IAUC 7597). We have previously argued that they might not be genuine supernovae, but rather superoutbursts of luminous variable stars (e.g., Filippenko et al. 1995, A.J. 110, 2261; Van Dyk et al. 2000, PASP 112, 1532). SN 2002kg provides strong evidence for this hypothesis; at a distance modulus of 27.6 mag (Freedman and Madore 1988, Ap.J. 332, L63), its absolute magnitude is only ∼ −9. However, these objects are still classified as supernovae, at least until they are convincingly shown not to be stellar explosions.”

SUPERNOVA 2002hz IN UGC 12044 Filippenko and Chornock add that inspection of CCD spectra, ob-

tained as above on Jan. 7 UT, shows that SN 2002hz (IAUC 8017) is of type Ib, roughly 2 months past maximum brightness and starting to enter the nebular phase. He I lines are clearly present, although their strength is in most cases difficult to judge owing to blending with other lines.

2003 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green